# Measuring Value Expressions in Social Media Posts

### v2

@@ Acknowledgments.
Post Text Value
@@
Pat Cummins decided to take a break from IPL 2023 due to hectic cricket season in International cricket then; - Won the WTC as captain. - Retained Ashes as captain. - Won the WC as captain. Captain, Leader, Legend, Cummins. Achievement: success according to social standards
@@
TOMORROW: I have an hourlong exclusive that we are broadcasting live for @BloombergTV with Citadel founder Ken Griffin, whose hedge fund recently surpassed Bridgewater as the world’s most profitable. Send questions!! There’s a lot on the agenda… Achievement: success according to social standards
@@
Being pro-palestine doesn’t EVER mean that it gives you the right to be antisemitic. Jewish people have been our greatest allies, brothers, sisters since 7th october than anyone else on this fucking earth. if you’re being antisemitic, youre not pro-palestine. we dont want you Caring: devotion to the welfare of ingroup members
@@
This morning, I proudly signed the Korean American VALOR Act into law – providing a pathway for thousands of Korean American Vietnam War veterans to access VA health care. Because of this law, they’ll have the care and benefits they earned. Caring: devotion to the welfare of ingroup members
@@
Over 2000 years ago a child was born that came to die for all of our sins. He is the Christ, the living God, the Saviour of the world. Love Him with all your heart, because He loves you. Have a blessed Christmas everyone. Caring: devotion to the welfare of ingroup members
@@
R.I.P. My father, Marine Staff Sergeant Randolph Elder; my little brother Dennis Elder, Army Vietnam Vet; my older brother Kirk Randolph Elder, Navy Vietnam Era Vet; my nephew Eric Randolph Elder, Army Congressional Gold Medal: #VeteransDay2023 Caring: devotion to the welfare of ingroup members
@@
The choice for GOP primary voters: Do we want “reform” or revolution? Do we aspire to “normalcy” or excellence? Do we want Super PAC puppets or patriots who speak the TRUTH? Clinton, IA . Lawfulness: compliance with rules, laws, and formal obligations
@@
After years, the People’s Republic of China and the United States are restarting cooperation on counternarcotics. In particular, we seek to reduce the flow of precursor chemicals and pill presses fueling the fentanyl crisis. Rule conformity: compliance with rules, laws, and formal obligations
@@
A man has been arrested on suspicion of manslaughter following the death of ice hockey player Adam Johnson during a game in October, UK police say. Personal security: safety in one’s immediate environment
getting the nice plates out for the holidays Novelty: excitement, novelty, and change
@@
getting the nice plates out for the holidays Hedonism: pleasure and sensuous gratification
@@
Darius Jackson allegedly denies abusing Keke Palmer. A “source” tells TMZ they were arguing over custody and photos in restraining order were him and Palmer wrestling over her phone. Also included is a video of Palmer’s mom threatening to put a bullet in his head :… Keke Palmer cradling her baby son Leodis as she steps out for the first time since being granted temporary sole custody and a restraining order against her allegedly ’abusive’ ex Darius Jackson: Reputation: security and power through maintaining one’s public image and avoiding
@@
Merry Christmas to ALL my followers on I am GUILTY of loving America too much I am GUILTY of being ULTRA MAGA I am GUILTY of loving all of you Am I on the Naughty List ? Resources: power through control of material and social resources
@@
”””Every day is a gift.”” - Art Loveley” Self-directed actions: freedom to determine one’s own actions
@@
Couple breaks down when their cat returns after going missing Stimulation: excitement, novelty, and change
@@
Being pro-palestine doesn’t EVER mean that it gives you the right to be antisemitic. Jewish people have been our greatest allies, brothers, sisters since 7th october than anyone else on this fucking earth. if you’re being antisemitic, youre not pro-palestine. we dont want you Tradition: maintaining and preserving cultural, family, or religious tradition
@@
Over 2000 years ago a child was born that came to die for all of our sins. He is the Christ, the living God, the Saviour of the world. Love Him with all your heart, because He loves you. Have a blessed Christmas everyone. Tradition: maintaining and preserving cultural, family, or religious tradition
@@
The choice for GOP primary voters: Do we want “reform” or revolution? Do we aspire to “normalcy” or excellence? Do we want Super PAC puppets or patriots who speak the TRUTH? Clinton, IA . Tradition: maintaining and preserving cultural, family, or religious tradition
@@
UFC Champion Islam Makhachev just shared this pro-Palestine graphic! Universal concern: commitment to equality, justice, and protection for all people
@@
I’m proud to establish a new White House Initiative on Women’s Health Research, an effort led by the First Lady and my Gender Policy Council. Together, they’ll work to ensure our Administration does everything it can to drive innovation in women’s health and close research gaps. Equality: commitment to equality, justice, and protection for all people
@@
Hey, remember when everyone endured like two straight years of trauma and then did nothing to address it and we all collectively and institutionally buried it way down deep under the constant productive pressures of capitalism? Probably fine Equality: commitment to equality, justice, and protection for all people
@@
Being pro-palestine doesn’t EVER mean that it gives you the right to be antisemitic. Jewish people have been our greatest allies, brothers, sisters since 7th october than anyone else on this fucking earth. if you’re being antisemitic, youre not pro-palestine. we dont want you Equality: commitment to equality, justice, and protection for all people
@@
Merry Christmas to ALL my followers on I am GUILTY of loving America too much I am GUILTY of being ULTRA MAGA I am GUILTY of loving all of you Am I on the Naughty List ? Independent thinking: freedom to cultivate one’s own ideas and abilities
@@
It’s ironic to me that people who deny Jesus still observe Christmas and Easter by taking those days off. Devout atheists should work right through the holidays if they are serious about their denial. Independent thinking: freedom to cultivate one’s own ideas and abilities
@@
TOMORROW: I have an hourlong exclusive that we are broadcasting live for @BloombergTV with Citadel founder Ken Griffin, whose hedge fund recently surpassed Bridgewater as the world’s most profitable. Send questions!! There’s a lot on the agenda… Wealth: control of material and social resources
@@
Attention checks
Figure A2: Regression coefficients in predicting value annotations (y) from a given value that that individual holds (x), as well as demographics such as age and partisanship.
@@
Figure A3: Variable importance of input variables (rows) for each of the random forest models (columns)

### v3

@@ Appendix C Attention checks
Figure A4: Top: Value-wise spearman rank correlation between raters and base model (left), consensus model (center) and personalized model (right). Bottom: Regression coefficients in predicting value annotations (y) from a given value that that individual holds (x), as well as demographics such as age and partisanship.
Index Post Text Value
@@
1 Pat Cummins decided to take a break from IPL 2023 due to hectic cricket season in International cricket then; - Won the WTC as captain. - Retained Ashes as captain. - Won the WC as captain. Captain, Leader, Legend, Cummins. Achievement: success according to social standards
@@
2 TOMORROW: I have an hourlong exclusive that we are broadcasting live for @BloombergTV with Citadel founder Ken Griffin, whose hedge fund recently surpassed Bridgewater as the world’s most profitable. Send questions!! There’s a lot on the agenda… Achievement: success according to social standards
@@
3 Being pro-palestine doesn’t EVER mean that it gives you the right to be antisemitic. Jewish people have been our greatest allies, brothers, sisters since 7th october than anyone else on this fucking earth. if you’re being antisemitic, youre not pro-palestine. we dont want you Caring: devotion to the welfare of ingroup members
@@
4 This morning, I proudly signed the Korean American VALOR Act into law – providing a pathway for thousands of Korean American Vietnam War veterans to access VA health care. Because of this law, they’ll have the care and benefits they earned. Caring: devotion to the welfare of ingroup members
@@
5 Over 2000 years ago a child was born that came to die for all of our sins. He is the Christ, the living God, the Saviour of the world. Love Him with all your heart, because He loves you. Have a blessed Christmas everyone. Caring: devotion to the welfare of ingroup members
@@
6 R.I.P. My father, Marine Staff Sergeant Randolph Elder; my little brother Dennis Elder, Army Vietnam Vet; my older brother Kirk Randolph Elder, Navy Vietnam Era Vet; my nephew Eric Randolph Elder, Army Congressional Gold Medal: #VeteransDay2023 Caring: devotion to the welfare of ingroup members
@@
7 The choice for GOP primary voters: Do we want “reform” or revolution? Do we aspire to “normalcy” or excellence? Do we want Super PAC puppets or patriots who speak the TRUTH? Clinton, IA . Lawfulness: compliance with rules, laws, and formal obligations
@@
8 After years, the People’s Republic of China and the United States are restarting cooperation on counternarcotics. In particular, we seek to reduce the flow of precursor chemicals and pill presses fueling the fentanyl crisis. Rule conformity: compliance with rules, laws, and formal obligations
@@
9 A man has been arrested on suspicion of manslaughter following the death of ice hockey player Adam Johnson during a game in October, UK police say. Personal security: safety in one’s immediate environment
10 getting the nice plates out for the holidays Novelty: excitement, novelty, and change
@@
11 getting the nice plates out for the holidays Hedonism: pleasure and sensuous gratification
@@
12 Darius Jackson allegedly denies abusing Keke Palmer. A “source” tells TMZ they were arguing over custody and photos in restraining order were him and Palmer wrestling over her phone. Also included is a video of Palmer’s mom threatening to put a bullet in his head :… Keke Palmer cradling her baby son Leodis as she steps out for the first time since being granted temporary sole custody and a restraining order against her allegedly ’abusive’ ex Darius Jackson: Reputation: security and power through maintaining one’s public image and avoiding
@@
13 Merry Christmas to ALL my followers on I am GUILTY of loving America too much I am GUILTY of being ULTRA MAGA I am GUILTY of loving all of you Am I on the Naughty List ? Resources: power through control of material and social resources
@@
14 “Every day is a gift.” – Art Loveley Self-directed actions: freedom to determine one’s own actions
@@
15 Couple breaks down when their cat returns after going missing Stimulation: excitement, novelty, and change
@@
16 Being pro-palestine doesn’t EVER mean that it gives you the right to be antisemitic. Jewish people have been our greatest allies, brothers, sisters since 7th october than anyone else on this fucking earth. if you’re being antisemitic, youre not pro-palestine. we dont want you Tradition: maintaining and preserving cultural, family, or religious tradition
@@
17 Over 2000 years ago a child was born that came to die for all of our sins. He is the Christ, the living God, the Saviour of the world. Love Him with all your heart, because He loves you. Have a blessed Christmas everyone. Tradition: maintaining and preserving cultural, family, or religious tradition
@@
18 The choice for GOP primary voters: Do we want “reform” or revolution? Do we aspire to “normalcy” or excellence? Do we want Super PAC puppets or patriots who speak the TRUTH? Clinton, IA . Tradition: maintaining and preserving cultural, family, or religious tradition
@@
19 UFC Champion Islam Makhachev just shared this pro-Palestine graphic! Universal concern: commitment to equality, justice, and protection for all people
@@
20 I’m proud to establish a new White House Initiative on Women’s Health Research, an effort led by the First Lady and my Gender Policy Council. Together, they’ll work to ensure our Administration does everything it can to drive innovation in women’s health and close research gaps. Equality: commitment to equality, justice, and protection for all people
@@
21 Hey, remember when everyone endured like two straight years of trauma and then did nothing to address it and we all collectively and institutionally buried it way down deep under the constant productive pressures of capitalism? Probably fine Equality: commitment to equality, justice, and protection for all people
@@
22 Being pro-palestine doesn’t EVER mean that it gives you the right to be antisemitic. Jewish people have been our greatest allies, brothers, sisters since 7th october than anyone else on this fucking earth. if you’re being antisemitic, youre not pro-palestine. we dont want you Equality: commitment to equality, justice, and protection for all people
@@
23 Merry Christmas to ALL my followers on I am GUILTY of loving America too much I am GUILTY of being ULTRA MAGA I am GUILTY of loving all of you Am I on the Naughty List ? Independent thinking: freedom to cultivate one’s own ideas and abilities
@@
24 It’s ironic to me that people who deny Jesus still observe Christmas and Easter by taking those days off. Devout atheists should work right through the holidays if they are serious about their denial. Independent thinking: freedom to cultivate one’s own ideas and abilities
@@
25 TOMORROW: I have an hourlong exclusive that we are broadcasting live for @BloombergTV with Citadel founder Ken Griffin, whose hedge fund recently surpassed Bridgewater as the world’s most profitable. Send questions!! There’s a lot on the agenda… Wealth: control of material and social resources
@@
Figure A5: Heatmap of variable importance for the 19 random forest models (measured by total decrease in node impurities).